OpenStudy (anonymous):

(3x10^-2)(2.1x10^5)... (^are exponents) Can I combine the 3x2.1 and the exponents together first or do I first need to simplify each term?

OpenStudy (anonymous):

multiply \(3\times 2.1\) and add the exponents

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3 \times 10^{-2} \times 2.1 \times 10^5}\) Don't get intimidated by those parentheses.

OpenStudy (anonymous):

Okay - that was my instinct - thank you thank you!

Parth (parthkohli):

Heh you're welcome!